# Recursive Entropy in Thermodynamics: Establishing the Statistical-Physics Basis of the Zentropy Approach

### v1

@@ 5 Derivation of the partition function
where the energy of the thermodynamic system is the expectation value of the fully specified configurations
$E=\langle E_{x}\rangle=\sum_{x}p_{x}E_{x}$ (43)
@@
For the case where partially specified configurations form a partition, as we did in the illustrative example, by the law of total expectation, we know that $\langle E_{x}\rangle=\langle E_{k}\rangle$, and the energy of the system can also be expressed as the expectation value over the partially specified configurations.
$E=\sum_{k}p_{k}E_{k}$ (44)
As each configuration is now treated as an ensemble, we would expect $F_{k}$ to take the place of $E^{k}$ in the standard approach. Indeed, this is what we find. Substituting equations Eq. (44) and Eq. (37) into the definition of Helmholtz energy gives
$F=\sum_{k}p_{k}E_{k}+T\left(\sum_{k}p_{k}\ln p_{k}-\sum_{k}p_{k}S_{k}\right)$ (45)
Then by the definition of Helmholtz energy Def. 5, $F_{k}=E_{k}-TS_{k}$, and Eq. (45) simplies to
$F=\sum_{k}p_{k}F_{k}+T\sum_{k}p_{k}\ln p_{k}$ (46)
With the thermodynamic system energy and Helmholtz energies now given in terms of $E_{k}$ and $S_{k}$ we can now derive the partition function in terms of $E_{k}$ and $S_{k}$. Following the standard derivation of the partition function, we maximize entropy $S=-\sum_{k}p_{k}\ln p_{k}+\sum p_{k}S_{k}$ under the constraints that the sum of the probabilities is equal to one
$\sum_{k}p_{k}=1$ (47)
and the internal energy is a constant
$U=\langle E\rangle=\sum_{k}p_{k}E_{k}$ (48)
The Lagrange function is
$\mathcal{L}=\left(-\sum_{k}p_{k}\ln p_{k}+\sum p_{k}S_{k}\right)+\alpha\left(1-\sum_{k}p_{k}\right)+\beta\left(U-\sum_{k}p_{k}E_{k}\right)$ (49)
To extremize entropy we set the variation of the Lagrange function to zero
@@
$=\sum_{k}\left[\delta\left(-p_{k}\ln p_{k}+p_{k}S_{k}\right)-\delta\left(\alpha p_{k}\right)-\delta\left(\beta E_{i}p_{i}\right)\right]$ (52)
$=\sum_{k}\left[\frac{\partial}{\partial p_{k}}\left(-p_{k}\ln p_{k}\right)\delta p_{k}+\frac{\partial}{\partial p_{k}}\left(p_{k}S_{k}\right)\delta p_{k}-\frac{\partial}{\partial p_{k}}\left(\alpha p_{k}\right)\delta p_{k}-\frac{\partial}{\partial p_{k}}\left(\beta E_{i}p_{i}\right)\delta p_{k}\right]$ (53)
$=\sum_{k}\left[-\ln p_{k}-1+S_{k}-\alpha-\beta E_{i}\right]\delta p_{k}$ (54)
relying on the fact that entropy is symmetric with respect to $p_{k}$, for any variation
$0=-\ln p_{k}-1+S_{k}-\alpha-\beta E_{k}$ (55)
@@
$p_{k}=e^{\left(-1-\alpha-\beta E_{k}+S_{k}\right)}$ (56)
applying our probability constraint Eq.(47)
@@
$\frac{1}{e^{-(1+\alpha)}}=\sum_{k}e^{\left(-\beta E_{k}+S_{k}\right)}$ (61)
@@
While we have derived the partition function in terms of $E_{k}$ and $S_{k}$, in order to write it in terms of $F_{k}$, we assume here that $\beta$ is still inverse temperature then prove it subsequently.
$Z$ $=\sum_{k}e^{\left(-\frac{1}{T}E_{k}+S_{k}\right)}$ (64)
$=\sum_{k}e^{\left(-\frac{1}{T}(E_{k}-TS_{k})\right)}$ (65)
$=\sum_{k}e^{\left(-\frac{F_{k}}{T}\right)}=\sum_{k}e^{\left(-\beta F_{k}\right)}$ (66)
To confirm that the identity of $\beta$ is inverse temperature and is unaffected by the inclusion of intra-configurational entropy, we write $p_{k}$ in terms of the partition function in order to write entropy in terms of the partition function and finally apply the definition of temperature. Substituting Eq. (63) into eq. (56),
$p_{k}=\frac{1}{Z}e^{\left(-\beta E_{k}+S_{k}\right)}$ (68)
Substituting into the definition of the logarithmic portion of entropy
@@
$=\beta E+\ln{Z}$ (75)
Differentiating with respect to energy gives
$\frac{\partial S}{\partial E}_{V,N}=\beta$ (77)
and by the definition of temperature $T=\frac{\partial E}{\partial S}_{V,N}$,
$\beta=\frac{1}{T}$ (78)
@@
Note that the definition of the partition function eq. (62) and the identity of $\beta$ are the same as in the standard thermodynamic approach, the derived relation between the partition function and $\beta$ is not. The difference being that $E_{x}$ is replaced by $F_{k}$, as expected.
@@ 6 Usefulness of zentropy
Why do we need recursive entropy when the definition is sufficient? Why group pure states? The answer is that recursive entropy is a generalization of the definition that provides a natural and efficiently computable approach to commonly encountered problems. Namely, problems where states of the system may be partitioned into basins. This partitioning into basins explains emergent pheonomena. Thus a framework, such as zentropy, that incorporates this partitioning of states, naturally predicts emergent phenomena.
@@
The recursive property of entropy (Eq. 1) is a generalization of the definition. If the set of states are partitioned so finely that each $k$ already specifies a pure state (the number of $k$ is maximal), then $S_{k}$ is equal to zero for all $k$, only the first term $k_{B}\sum_{k}p_{k}\ln p_{k}$ remains, $k=x$, and we recover the definition of entropy (Eq. 2). Keeping the intra-configurational term therefore just allows us to choose configurations that contain additional unresolved mixed-up-ness. Recursive entropy is thus a generalization and reduces to the definition when no hidden uncertainty remains in each configuration ($S_{k}=0)$.
@@
Conversely, at the opposite extreme, we could choose our set of states to be as coarsely partitioned as possible so that there is only one configuration (The number of $k$ is minimal), in which case, the configurational entropy term is zero $k_{B}\sum_{k}p_{k}\ln p_{k}=0$ and Eq. 1 becomes trivial $S=\sum_{k}p_{k}S_{k}=S_{k}$. However, it is often best to choose a partitioning that is somewhere inbetween. All this is to say that by including the intra-configurational entropy term $\sum_{k}p_{k}S_{k}$—using the recursive property of entropy— allows us to choose a definition of configuration that is most useful. Or in other words, to define the configurations of the system at the most useful scale.
@@
A useful way to group pure states is in the same way as they are partitioned in nature, that is, to partition pure states in accordance with the basins of the energy landscape. This way, it is natural to consider the system’s behavior given a specifed basin/configuration. This can be used to simplify calculations by applying approximations hierarchically, as the pure states in each basin often have some shared properties. For example, consider a magnetic crystal, where each configuration is a given magnetic ordering. The remaining uncertainty is then in the phonon and electron states given that magnetic ordering.
$S_{k}=S_{k,\text{el}}+S_{k,\text{vib}}$ (79)
@@
where $S_{k}$ is the entropy of a given configuration/ magnetic ordering, $S_{\text{el}}$ is entropy due to electron excitations, and $S_{\text{vib}}$ due to lattice vibrations. Practically, this may be calculated using density functional theory. The additivity of $S_{\text{el}}$ and $S_{\text{vib}}$ assumes that the phonon density of states and the electron density of states are independent for each given configuration. This assumption facilitates the calculations. Otherwise, they are subadditive, and an additional relation between $S_{\text{el}}$ and $S_{\text{vib}}$ is required to determine the $S_{k}$ exactly. However, such a calculation is difficult and often unnecessary.
@@
Here we clarify some terminology. There are two different uses of the term ”coarse-graining”. To clarify, one meaning refers to the partitioning or, more generally, grouping of states, and the other refers to the incomplete sampling of configurations. While the term ”coarse-graing” is often used for both, The two meanings are very different. Notably, incomplete sampling is approximate, whereas grouping of states and calculating entropy with the recursive property is exact.
@@
We now turn our attention to the incomplete sampling of configurations. While approximate, it is often not feasible to calculate all possible configurations of a system as required by the partition function. Instead, a representative sample of configurations may be used. For example, this may be achieved by calculating all the configurations of a finite cell rather than the infinite lattice, or using Monte Carlo or other methods to sample only configurations of sufficient probability. By grouping pure states, we can more efficiently sample or approximate the pure states. For example, by partitioning pure states by spin configuration, we can apply the harmonic approximation to each configuration to gather the phonon density of states.
Including both configurational and intra-configurational terms in the equation for entropy, allows defining configurations on a scale between the macroscopic thermodynamic state and quantum pure states. As mentioned earlier, for partitions, when the number of configurations $k$ is maximal, the configurations are defined on the scale of quantum pure states, and when the number of configurations is minimal, the configuration is the macroscopic thermodynamic state. By choosing the partition such that the number of configurations is between the two extremes, we build a bridge between scales.
@@
This bridge can be useful in the description of emergent phenomena. While there are rules that govern the quantum pure states (quantum mechanics), one could form new rules that govern the configurations formed by the grouping. Thus, a rule set for the quantum pure states, and an emergent rule set for the configurations govern the same system. This approach works particularly well for configurations wherein the quantum pure states (or more gennerally subconfigurations) are strongly lumpable [Rosas2024]. That is, the quantum pure states of one group have a defining property that is clearly distinct from quantum pure states of another configuration.
In summary, the recursive property of entropy provides a flexible and exact framework for partitioning thermodynamic systems at physically meaningful scales. By grouping pure states according to natural basins in the energy landscape, we can apply hierarchical approximations that exploit shared properties within each configuration, simplifying otherwise intractable calculations. The multiscale nature of the approach naturally bridges microscopic and macroscopic descriptions. This bridging capability makes zentropy particularly powerful for describing emergent phenomena, where different rule sets govern behavior at different scales, provided the configurations are sufficiently distinct or ”strongly lumpable.” The result is a computationally practical approach that maintains rigorous statistical-mechanical foundations while enabling quantitative predictions across scales. Here, an analogy can be drawn between zentropy and Kohn–Sham DFT, which provides a practical approach to quantum mechanics by replacing an intractable interacting many-electron problem with a non-interacting system in an effective potential that preserves the exact density.
@@ 7 Conclusion
We have established a clear statistical-mechanical foundation for the zentropy approach by proving the recursive property of entropy as a hierarchical form of the chain rule and showing its equivalence to the standard definition when configurations reduce to pure states. Building on this foundation, we derived a grouped form of the Helmholtz energy and obtained a partition function expressed in terms of configuration Helmholtz energies. This zentropy approach separates configurational entropy from intra-configurational contributions and provides a principled, exact route to multiscale modeling, by aligning configurations with natural basins in the energy landscape
Practically, these results position zentropy as a computationally efficient bridge between microscopic states and macroscopic behavior, clarifying when and how emergent phenomena can be captured by a reduced description. We anticipate that adopting the recursive property in thermodynamics and condensed-matter calculations will streamline predictive workflows, particularly in DFT and when configurations are strongly lumpable.
@@ 8 Acknowledgments
The present work is funded by the Department of Energy (DOE) through Grant No. DE-SC0023185

### v2

@@ 5 Helmholtz energy and the Partition Function
where the internal energy $U$ of the thermodynamic system is the expectation value of the energy of the fully-constrained configuration:
$U=\langle E_{x}\rangle=\sum_{x}p_{x}E_{x}.$ (46)
@@
For the case where the partially-constrained configurations form a partition, as we did in the illustrative example in Section 3, the law of total expectation applies so $\langle E_{x}\rangle=\langle E_{k}\rangle$, and the energy of the system can also be expressed as the expectation value over the partially-constrained configurations:
$U=\sum_{k}p_{k}E_{k}.$ (47)
Naively, we might expect that $F=\langle F\rangle=\sum_{k}p_{k}F_{k}$. However, substituting Eq. (47) and Eq. (37) into the definition of Helmholtz energy (Eq. (45)) gives
$F=\sum_{k}p_{k}E_{k}+T\left(\sum_{k}p_{k}\ln p_{k}-\sum_{k}p_{k}S_{k}\right).$ (48)
For convenience, we define the quantity $F_{k}=E_{k}-TS_{k}$, and Eq. (48) simplifies to
$\boxed{F=\sum_{k}p_{k}F_{k}+T\sum_{k}p_{k}\ln p_{k}.}$ (49)
Indeed, for partially-constrained configurations that form a partition, the Helmholtz energy $F_{k}$ replaces the energy $E_{x}$ of fully-constrained configurations.
Next, we derive the partition function by maximizing entropy in its recursive form (Eq. (37)) under the constraints that the sum of the probabilities is equal to one,
$\sum_{k}p_{k}=1,$ (50)
and that the expectation value of the energy is a constant,
$U=\sum_{k}p_{k}E_{k}.$ (51)
Following the method of Lagrange multipliers in the calculus of variations, the Lagrange function is
$\mathcal{L}=\left(-\sum_{k}p_{k}\ln p_{k}+\sum p_{k}S_{k}\right)+\alpha\left(1-\sum_{k}p_{k}\right)+\beta\left(U-\sum_{k}p_{k}E_{k}\right).$ (52)
To extremize entropy, we set the variation of the Lagrange function to zero and simplify:
@@
$=\sum_{k}\left[\delta\left(-p_{k}\ln p_{k}+p_{k}S_{k}\right)-\delta\left(\alpha p_{k}\right)-\delta\left(\beta E_{k}p_{k}\right)\right]$ (55)
$=\sum_{k}\left[\frac{\partial}{\partial p_{k}}\left(-p_{k}\ln p_{k}\right)\delta p_{k}+\frac{\partial}{\partial p_{k}}\left(p_{k}S_{k}\right)\delta p_{k}-\frac{\partial}{\partial p_{k}}\left(\alpha p_{k}\right)\delta p_{k}-\frac{\partial}{\partial p_{k}}\left(\beta E_{k}p_{k}\right)\delta p_{k}\right]$ (56)
$=\sum_{k}\left[-\ln p_{k}-1+S_{k}-\alpha-\beta E_{k}\right]\delta p_{k}.$ (57)
Since the variations are independent, the coefficient of each $\delta p_{k}$ must vanish. Thus, for every $k$,
$0=-\ln p_{k}-1+S_{k}-\alpha-\beta E_{k}.$ (58)
@@
$p_{k}=e^{\left(-1-\alpha-\beta E_{k}+S_{k}\right)}.$ (59)
Applying the probability constraint (Eq. (50)),
@@
$\frac{1}{e^{-(1+\alpha)}}=\sum_{k}e^{\left(-\beta E_{k}+S_{k}\right)}.$ (64)
@@
Assuming, that $\beta$ is inverse temperature, we can write the partition function in terms of $F_{k}$,
$\boxed{Z=\sum_{k}e^{\left(-\frac{F_{k}}{T}\right)}=\sum_{k}e^{\left(-\beta F_{k}\right)}.}$ (67)
To confirm that the identity of $\beta$ is indeed inverse temperature and is unaffected by the inclusion of intra-configurational entropy, we write $p_{k}$ and entropy $S$ in terms of the partition function and then apply the definition of temperature. Substituting Eq. (65) into Eq. (59),
$p_{k}=\frac{1}{Z}e^{\left(-\beta E_{k}+S_{k}\right)}.$ (68)
Next, we rewrite the recursive property (Eq. (37)) using Eq. (68) to evaluate $\ln p_{k}$ (leaving the prefactor and other instances of $p_{k}$ unchanged:
@@
$=\beta U+\ln{Z}.$ (75)
Notice that the $S_{k}$ terms cancel out in Eq. (71). Differentiating with respect to the internal energy gives
$\left(\frac{\partial S}{\partial U}\right)_{V,N}=\beta,$ (76)
and by the definition of temperature $T=\left(\frac{\partial U}{\partial S}\right)_{V,N}$,
$\beta=\frac{1}{T},$ (77)
@@
Note that while the definition of the partition function (Eq. (65)) and the identity of $\beta$ (Eq. (77)) are the same as in the standard thermodynamic approach, the derived relation between the partition function and $\beta$ (Eq. (67)) is not. The difference being that energy is replaced by Helmholtz energy, as expected.
@@ 6 Usefulness of Zentropy
Why do we need recursive entropy (Eq. (2)) when we have the standard definition (Eq. (1))? Why group microstates at all? The answer is that recursive entropy is a generalization of the standard definition that provides a natural and computationally efficient approach for commonly encountered problems. Namely, problems where microstates of the system may be partitioned into basins. This partitioning into basins explains emergent phenomena. Thus, a framework, such as zentropy, that incorporates a hierarchical partitioning of microstates, naturally predicts emergent phenomena.
@@
The recursive property of entropy reduces to the standard definition in the limit of fully-constrained configurations. If the set of microstates is partitioned so finely that each $k$ already specifies a fully-constrained configuration (the number of $k$ is maximal), then $S_{k}$ is equal to zero for all $k$, only the first term $k_{B}\sum_{k}p_{k}\ln p_{k}$ remains. Identifying $k=x$, we recover the usual definition of entropy.
@@
Conversely, at the opposite extreme, we could choose our set of microstates to be as coarsely partitioned as possible, so that there is only one configuration (the number of $k$ is minimal), in which case, the configurational entropy term vanishes ($k_{B}\sum_{k}p_{k}\ln p_{k}=0$) and Eq. (2) becomes trivial $S=\sum_{k}p_{k}S_{k}=S_{k}$. In practice, however, the most useful description typically lies between these two extremes. All this is to say that including the intra-configurational entropy term $\sum_{k}p_{k}S_{k}$ allows us to choose a definition of configuration that is best suited to the problem at hand. Or, in other words, to define the configurations of the system at the most useful scale. We emphasize that the recursive property is not a statement about the intrinsic nature of a system; rather, its usefulness lies in the flexibility it gives us in describing the system in terms of partially-constrained configurations.
@@
A useful way to group microstates is in the same way as they are partitioned in nature, that is, to partition microstates in accordance with the basins of the energy landscape. This way, it is natural to consider the system’s behavior given a specified basin/configuration. This can be used to simplify calculations by applying approximations hierarchically, as the microstates in each basin often have shared properties. For example, consider a magnetic crystal, where each configuration is a given magnetic ordering. The remaining uncertainty is then in the phonon and electron states given that magnetic ordering:
$S_{k}=S_{k,\text{el}}+S_{k,\text{vib}},$ (78)
@@
where $S_{k}$ is the entropy of a given configuration/magnetic ordering, $S_{\text{el}}$ is entropy due to electron excitations, and $S_{\text{vib}}$ due to lattice vibrations. Practically, this may be calculated using density functional theory. The additivity of $S_{\text{el}}$ and $S_{\text{vib}}$ assumes that the phonon density of states and the electron density of states are independent for each given configuration. This assumption facilitates the calculations. Otherwise, they are subadditive, and an additional relation between $S_{k,\text{el}}$ and $S_{k,\text{vib}}$ is required to determine $S_{k}$ exactly. However, such a calculation is difficult and often unnecessary [Shang2010b, Wang2004].
@@
Here, we clarify some terminology. There are two different uses of the term “coarse-graining”. The first, which we have already dealt with, refers to the partitioning or, more generally, grouping of microstates, while the other refers to the incomplete sampling of configurations. Although the term “coarse-graining” may be used for both [Bengtsson2017, Kmiecik2016], the two meanings are fundamentally different. Notably, incomplete sampling is approximate, whereas grouping of microstates and calculating entropy with the recursive property is exact.
@@
We now turn our attention to the incomplete sampling of configurations. Although approximate, it is often infeasible to calculate all possible configurations of a system, as required by the partition function. Instead, a representative sample of configurations may be used. For example, this may be achieved by calculating all the configurations of a finite cell rather than the infinite lattice, or by employing Monte Carlo or related methods to sample only configurations with sufficiently high probability.
@@
Including both configurational and intra-configurational terms in the equation for entropy allows defining configurations on a scale between the macroscopic thermodynamic state and fully-constrained microscopic pure states. Thus, we build a bridge between scales, as illustrated in Figure 1b, which can be useful in the description of emergent phenomena. While there are rules that govern the fully-constrained configurations (quantum mechanics), one can define new, emergent rules that govern the grouped configurations. In this way, a system can be described both by the rule set for fully-constrained configurations and by an emergent rule set for the grouped configurations. This approach works particularly well for configurations wherein the subconfigurations are strongly lumpable [Rosas2024]. That is, the subconfigurations of one configuration have a defining property that is clearly distinct from the subconfigurations of another configuration.
Ultimately, the utility of zentropy lies in the ability to describe systems accurately despite the practical necessity of incomplete sampling. By providing a mathematical structure that adapts to the natural basins of the energy landscape, the framework renders microscopic complexity tractable on macroscopic scales. This flexibility streamlines computation and illuminates the causal links between fundamental interactions and emergent macroscopic behavior.
@@ 7 Conclusion
In summary, the recursive property of entropy provides a flexible and exact framework for partitioning thermodynamic systems at physically meaningful scales. By grouping microstates according to natural basins in the energy landscape, we can apply hierarchical approximations within each configuration, exploiting the shared properties of the microstates and thereby simplifying otherwise intractable calculations. The multiscale nature of the approach naturally bridges microscopic and macroscopic descriptions, which makes zentropy particularly powerful for describing emergent phenomena, where different rule sets govern behavior at different scales, provided the subconfigurations are sufficiently distinct or “strongly lumpable.” The result is a computationally practical approach that maintains rigorous statistical-mechanical foundations while enabling quantitative predictions across scales. Here, an analogy can be drawn between zentropy and Kohn–Sham DFT, which provides a practical approach to quantum mechanics by replacing an intractable interacting many-electron problem with a non-interacting system in an effective potential that preserves the exact density.
We have established a clear statistical-mechanical foundation for the zentropy approach by proving the recursive property of entropy as a hierarchical form of the chain rule and showing its equivalence to the standard definition when configurations are fully constrained. Building on this foundation, we derived a grouped form of the Helmholtz energy and obtained a partition function expressed in terms of configuration Helmholtz energies. This zentropy approach separates configurational entropy from intra-configurational contributions and provides a principled, exact route to multiscale modeling, by aligning configurations with natural basins in the energy landscape.
Practically, these results position zentropy as a computationally efficient bridge between microscopic and macroscopic behavior, clarifying when and how emergent phenomena can be captured by a reduced description. We anticipate that adopting the recursive property in thermodynamics and condensed-matter calculations will streamline predictive workflows, particularly in DFT and when subconfigurations are strongly lumpable.
@@ 8 Acknowledgments
The present work is funded by the Department of Energy (DOE) through Grant No. DE-SC0023185.
CRediT author statement
Luke Allen Myers: Conceptualization, Methodology, Writing - Original Draft, Writing - Review & Editing, Visualization
Nigel Lee En Hew: Conceptualization, Writing - Review & Editing, Supervision
Shun-Li Shang: Writing - Review & Editing, Supervision, Funding acquisition
Zi-Kui Liu: Writing - Review & Editing, Supervision, Funding acquisition